# Superfast Low Rank Approximation

### v2

@@ 9 Numerical tests
In this section we cover our tests of dual superfast variants of Alg. 6.1. As in the tests of Part I we applied the standard normal distribution function randn of MATLAB to generate Gaussian matrices. The MATLAB function ”svd()” has been applied to calculate the $\epsilon$-rank for $\epsilon=10^{-6}$. The tests for Tables 9.19.3 have been performed on a 64-bit Windows machine with an Intel i5 dual-core 1.70 GHz processor by using custom programmed software in $C^{++}$ and compiled with LAPACK version 3.6.0 libraries.
@@ 9.1 Input matrices
Class II: (i) The matrices $M$ of the discretized single-layer Laplacian operator of [39, Sec. 7.1]: $S\sigma](x)=c\int_{\Gamma_{1}}\log{|x-y|}\sigma(y)dy,x\in\Gamma_{2}$, for two circles $\Gamma_{1}=C(0,1)$ and $\Gamma_{2}=C(0,2)$ on the complex plane. We arrived at the matrices $M=(m_{ij})_{i,j=1}^{n}$, $m_{i,j}=c\int_{\Gamma_{1,j}}\log|2\omega^{i}-y|dy$ for a constant $c$, M=1||M||=1 and the arc $\Gamma_{1,j}$ of $\Gamma_{1}$ defined by the angles in the range $\frac{2j\pi}{n},\frac{2(j+1)\pi}{n}]$.
(ii) The matrices that approximate the inverse of a large sparse matrix obtained from a finite-difference operator of [39, Sec. 7.2].
@@
9.2 Five families of Ultrasparse sketch matrices $H$
@@ 9.3 Test results
Tables 9.19.6 display the average relative error norm $\frac{\|M-\tilde{M}\|_{2}}{\|M-M_{nrank}\|_{2}}$ in our tests repeated 100 times for each class of input matrices and each size of an input matrix and sketch matrix $H$ for Alg. 6.1 or for each size of an input matrix and a pair of left-hand and right-hand sketch matrices $F$ and $H$ for Alg. 6.3.
In all our tests we applied the sketch matrices of the six families of the previous subsection.
Tables 9.19.3 display the average relative error norm for the output of Alg. 6.1; in our tests it ranged from about $10^{-3}$ to $10^{1}$. The numbers in parentheses in the first line of Tables 9.2 and 9.3 show the numerical rank of input matrices.
Tables 9.49.6 display the average relative error norm for the output of Alg. 6.3 applied to the same input matrices from classes I–III as in our experiments for Alg. 6.1.
In these tests we used $n\times\ell$ and $\ell\times m$ sketch matrices for $\ell=r+p$ and $k=c\ell$ for $c=1,2,3$ and random $p=1,2,\dots,21$.
@@ 10 Conclusions
[53] proposed and tested iterative refinement of LRA, and [28] has greatly advanced that study and incorporated the superfast algorithms of our present paper.
@@
[1] computed least squares approximation of multivariate functions (LSAMF) by means of reduction to the solution of Linear Least Squares Problem (LLSP) and further to C-A iterations for LRA but actually to any LRA algorithm, including our superfast ones.
@@ Appendix A Small families of hard inputs for superfast LRA
Appendix B LRA with dense and Abridged SRHT random sketch matrices
@@ B.1 The generalized Nyström LRA algorithm (briefly)
Choose the generalized Nyström randomized algorithm of [48] to represent random sketching LRA algorithms. It modifies the algorithms of [18, Thm. 4.3, display 1] and [70] to improve numerical stability of rank-$r$ approximation of a matrix $M\in\mathbb{C}^{m\times n}$ for $r\leq n\leq m$. It first generates two random matrices $F\in\mathbb{C}^{k\times m}$ and $H\in\mathbb{C}^{n\times\ell}$, for $r\leq k\leq\ell\leq n$, then computes the sketches $FM$ and $MH$, and finally computes rank-$r$ approximation $\{X,Y,Z\}$ of $M$.
In the best studied case $F$ and $H$ are Gaussian random matrices, filled with independent standard Gaussian (normal) random variables; then for $k=2r+1$ and $\ell=2k$ the expected Frobenius error norm of the error matrix MXYZM-XYZ is within a factor of 2 from optimal (not exceeding the optimal one for rank-$\frac{k}{2}$ approximation). The algorithm involves $O((k+\ell)mn)$ flops, $mn$ entries of $M$, and about $km+n\ell+k\ell$ other scalars, overwriting some entries of $M$.
@@
Remark B.1.
[Superfast LRA under the spectral norm.] The known algorithm supporting the same LRA error bounds under the spectral rather than Frobenius norm tend to be more involved (see [39, 25, 69]). E.g., the Power algorithm of [39] computes matrices $F(MM^{*})^{j}M$ and $(MM^{*})^{j}MH$, for random matrices $F$ and $H$, $j=0,1,\dots,t$, and $t$ of order of $\log(m+n)$. One can accelerate these computations by using sparse and/or structured matrices $F$ and $H$ but only for $j=0$ because multiplication by $MM^{*}$ destroys sparsity and structure.
@@ B.2 Abridged SRHT matrices
With sparse subspace embedding [9, 12, 14], [71, Sec. 3.3], [46, Sec. 9] one obtains significant acceleration but still does not yield superfast algorithms. According to [40], such acceleration tends to make the accuracy of the output LRAs somewhat less reliable, although [14] partly overcomes this problem for incoherent matrices. One can multiply a matrix by Subsampled Randomized Hadamard or Fourier Transform (SRHT or SRFT) dense matrices towards incoherence [14], but this step is not superfast. By applying recipe of the end of Sec. 1.4, we devise Abridged SRHT multipliers, recall Thm. 8.2 for semi-empirical support for the output accuracy of the resulting LRA, and recall [57] for earlier formal and empirical study of their application to LRA.
We proceed by means of abridging the classical recursive processes of the generation of $n\times n$ SRHT matrices, obtained from the $n\times n$ dense matrices $H_{n}$ of Walsh-Hadamard transform for $n=2^{t}$ (cf. [44, Sec. 3.1]). Namely, we write
@@
We also refer to $d$-abridged SRHT matrices as $d$-Abridged Scaled and Permuted Hadamard $d$-ASPH matrices and call them $d$-APH matrices if random scaling is omitted, that is, if the multiplier $D$, filled with $\pm 1$, is replaced with the identity matrix.
@@ Appendix C The spectral norms of a Gaussian matrix and its pseudo inverse
[Spectral norms of a Gaussian matrix. See [21, Thm. II.7].]
@@
See [10, Proof of Lemma 4.1] for claim (i), [39, Prop. 10.4 and Eqns. (10.3) and (10.4)] for claims (ii) and (iii), and [65, Thm. 3.3] for claim (iv). ∎

### v3

@@ 10 Numerical tests
Next we cover our tests of dual superfast variants of Algs. 7.1 and 7.3 specified in Remark 7.2. As in Sec. 6 we applied the standard normal distribution function randn of MATLAB to generate Gaussian matrices. We apply MATLAB function ”svd()” to calculate the $\epsilon$-rank for $\epsilon=10^{-6}$ and perform the tests for Tables 10.110.3 on a 64-bit Windows machine with an Intel i5 dual-core 1.70 GHz processor by using custom programmed software in $C^{++}$ and compiled with LAPACK version 3.6.0 libraries, as in Sec. 6.
@@ 10.1 Input matrices
Class II: (i) The matrices $M$ of the discretized single-layer Laplacian operator of [30, Sec. 7.1]: $S\sigma](x)=c\int_{\Gamma_{1}}\log{|x-y|}\sigma(y)dy,x\in\Gamma_{2}$, for two circles $\Gamma_{1}=C(0,1)$ and $\Gamma_{2}=C(0,2)$ on the complex plane. We arrived at the matrices $M=(m_{ij})_{i,j=1}^{n}$, $m_{i,j}=c\int_{\Gamma_{1,j}}\log|2\omega^{i}-y|dy$ for a constant $c$, M=1||M||=1 and the arc $\Gamma_{1,j}$ of $\Gamma_{1}$ defined by the angles in the range $\frac{2j\pi}{n},\frac{2(j+1)\pi}{n}]$.
(ii) The matrices that approximate the inverse of a large sparse matrix obtained from a finite-difference operator of [30, Sec. 7.2].
@@
10.2 Five families of Ultrasparse multipliers $H$
@@ 10.3 Test results
Tables 10.110.6 display the average relative error norm $\frac{\|M-\tilde{M}\|_{2}}{\|M-M_{nrank}\|_{2}}$ in our tests repeated 100 times for each class of input matrices and each size of an input matrix and multiplier $H$ for Alg. 7.1 or for each size of an input matrix and a pair of left-hand and right-hand multipliers $F$ and $H$ for Alg. 7.3.
In all our tests we applied the multipliers of the six families of the previous subsection.
Tables 10.110.3 display the average relative error norm for the output of Alg. 7.1; in our tests it ranged from about $10^{-3}$ to $10^{1}$. The numbers in parentheses in the first line of Tables 10.2 and 10.3 show the numerical rank of input matrices.
Tables 10.410.6 display the average relative error norm for the output of Alg. 7.3 applied to the same input matrices from classes I–III as in our experiments for Alg. 7.1.
In these tests we used $n\times\ell$ and $\ell\times m$ multipliers for $\ell=r+p$ and $k=c\ell$ for $c=1,2,3$ and random $p=1,2,\dots,21$.
@@ 11 Conclusions
Next we list some promising directions for further study of superfast LRA.
[44] proposed and tested iterative refinement of LRA, and [21] has greatly advanced that study incorporating the superfast algorithms that we studied here.
@@
Least squares approximation of multivariate functions (LSAMF) was reduced in [1] to the solution of Linear Least Squares Problem (LLSP) and further to LRA. Namely, [1] reduced LSAMF to LLSP, which we state as follows: for given $M\in\mathbb{C}^{m\times n}$ and ${\bf b}\in\mathbb{C}^{m\times 1}$ where $m>n$, compute or approximate a vector
[…]
Triangular inequality implies that for an LRA $\tilde{M}$ of $M$ we have
[…]
[1] showed that $(\tilde{M}-M){\bf y}||=0$ in the case of LSAMF problem, and so we can solve it superfast if LRA can be computed superfast, e.g., with our algorithms.
Indeed, suppose that our superfast LRA algorithm has computed $P$ and $S$ such that $\tilde{M}=PS\in\mathbb{R}^{m\times n}$, $P\in\mathbb{R}^{m\times r}$, $S\in\mathbb{R}^{r\times n}$ and $r\ll\min\{m,n\}$. Then compute a solution
[…]
@@ Appendix A Small families of hard inputs for superfast LRA
Appendix B LRA with dense and Abridged SRHT random multipliers
@@ B.1 The generalized Nyström LRA algorithm (briefly)
The generalized Nyström algorithms of [60, 39] specify Alg. 7.3 for random sketching LRA.131313[39] modifies the algorithms of [60] to improve numerical stability of rank-$r$ approximation of a matrix $M\in\mathbb{C}^{m\times n}$ for $r\leq n\leq m$, but in our tests we departed from [60] and accelerated it by using Abridging SRHT multipliers $F$ and $H$. In the best studied case the multipliers $F$ and $H$ are Gaussian random matrices, filled with independent standard Gaussian (normal) random variables; then for $k=2r+1$ and $\ell=2k$ the expected Frobenius error norm of the error matrix MXYZM-XYZ is within a factor of 2 from optimal (not exceeding the optimal one for rank-$\frac{k}{2}$ approximation). The algorithm involves $O((k+\ell)mn)$ flops, $mn$ entries of $M$, and about $km+n\ell+k\ell$ other scalars, overwriting some entries of $M$.
@@ B.2 Abridged SRHT matrices
With sparse subspace embedding [10, 12], [61, Sec. 3.3], [37, Sec. 9] one obtains significant acceleration but still does not yield superfast algorithms. According to [32], such acceleration tends to make the accuracy of the output LRAs somewhat less reliable, although [12] partly overcomes this problem for incoherent matrices. One can multiply a matrix by Subsampled Randomized Hadamard or Fourier Transform (SRHT or SRFT) dense matrices towards incoherence [12], but this step is not superfast. We make it superfast by applying Abridged SRHT multipliers, specified below. Thm. 9.1 provides semi-empirical support for the output accuracy of the resulting LRA. [48] covers earlier formal and empirical study of Abridged SRHT matrices and their application to LRA.
We proceed by means of abridging the classical recursive processes of the generation of SRHT matrices, obtained from the $n\times n$ dense matrices $H_{n}$ of Walsh-Hadamard transform for $n=2^{t}$ (cf. [35, Sec. 3.1]). Namely, we write
@@
We also refer to $d$-abridged SRHT matrices as $d$-Abridged Scaled and Permuted Hadamard $d$-ASPH matrices and call them $d$-APH matrices if random scaling is omitted, that is, if the multiplier $D$, filled with $\pm 1$, is replaced with the identity matrix, and then randomization only comes from random permutation matrix.
@@ Appendix C The spectral norms of a Gaussian matrix and its pseudo inverse
[Spectral norms of a Gaussian matrix. See [17, Thm. II.7].]
@@
See [8, Proof of Lemma 4.1] for claim (i), [30, Prop. 10.4 and Eqns. (10.3) and (10.4)] for claims (ii) and (iii), and [57, Thm. 3.3] for claim (iv). ∎
@@
Appendix D Randomized pre-processing of LRA
The following simple results from [47, Section 8.2] show that pre-processing with Gaussian multipliers $X$ and $Y$ transforms any matrix that admits $LRA$ into a perturbation of a factor-Gaussian matrix. We write $A\preceq B$ (resp. $A\succeq B$) to show that $A$ is statistically less (resp. greater) than or equal to $B$.
Theorem D.1.
Consider five integers $k$, $l$, $m$, $n$, and $r$ satisfying the bounds $r\leq k\leq m,~r\leq l\leq n$, an $m\times n$ well-conditioned matrix $M$ of rank $r$, $k\times m$ and $n\times l$ Gaussian matrices $G$ and $H$, respectively, and the norms $\nu_{p,q}$ and $\nu_{p,q}^{+}$ of Appendix C. Then
(i) $GM$ is a left factor-Gaussian matrix of expected rank $r$ such that
[…]
(ii) $MH$ is a right factor-Gaussian matrix of expected rank $r$ such that
[…]
(iii) $GMH$ is a two-sided factor-Gaussian matrix of expected rank $r$ such that
[…]
Remark D.1.
According to this theorem we can readily extend our results on $LRA$ of perturbed factor-Gaussian matrices to all matrices that admit $LRA$ and are pre-processed with Gaussian multipliers. We can perform such pre-processing superfast if we replace Gaussian multipliers with 3-abridged SRHT orthogonal multipliers; this replacement has little affected the output accuracy of LRA for a large class of matrices $M$ in the tests in [21].
Appendix E Proof of Thm. 2.1
Proof.
Now let $M=M^{\prime}=CUR$. Then clearly
Hence
It remains to deduce that
but in this case $G_{r}=G$, and so
Remark E.1.
$M=\begin{pmatrix}1&0&0\\ 0&\epsilon&0\\ 0&1&0\end{pmatrix}$,